using the elimination method how would you solve this:
2x+3y=-7
x-4y=-9

Answer by user_dude2008(1862)

2x+3y=-7
x-4y=-9


2x+3y=-7
-2x+8y=18

11y=11
y=1

x-4y=-9
x-4(1)=-9
x-4=-9
x=-5

x=-5, y=1